OpenStudy (anonymous):

Brad wants to buy flowers for his friend with $27. The daisies are $1 each and the roses are $3 each. He buys 3 more daisies than roses. Q: How much did the roses cost?

OpenStudy (bibby):

Th problem states that the roses are the variable (changing) factor We know that no matter how many roses he has, he'll always have 3 more daisies, right? In equation form. that's (3*x)+(3*1)

OpenStudy (bibby):

(we're multiplying by the prices). Should I go in further detail? 3x+3=27

OpenStudy (anonymous):

Yes Please

OpenStudy (bibby):

step 1: find the variable here it says: roses = x daisies = x+3 We couldn't set x+x+3 = $27 because the left hand side is the amount of roses not the cost

OpenStudy (anonymous):

Oh Okay

OpenStudy (bibby):

I meant amount of flowers (not just roses) However if we multiply by their costs we get: (x) times $3 - roses (x+3) times $1 -daisies 3x+x+3=27

OpenStudy (anonymous):

I dont understand what the answer would be.. :p

OpenStudy (bibby):

The answer isn't as important as the road we take to get there :D Do you know how to approach 3x+x+3=27 ?

OpenStudy (anonymous):

No I Do Not :p

OpenStudy (anonymous):

Explain Please....

OpenStudy (bibby):

okay let's see. as with any equations you want to combine like terms. That means adding and subtracting up the x's that you can 3x+x=4x So now we're left with 4x+3=27 next we try and get the x alone

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

Subract 4 from each side?

OpenStudy (bibby):

you can't do that as the 4 is multiplied to the x (did you mean 3?) That's why in general we'd subtract 3 first

OpenStudy (anonymous):

Oh Its 3!? Oh ! Okay

OpenStudy (bibby):

The answer? WE still have a few more steps lol if we subtract 3 on both sides we get 4x=24 THEN we can divide to get the x alone x=6 actually I think I made a mistake. give me a minute

OpenStudy (bibby):

He was supposed to have bought 8 roses. I'm sorry about this mixup. I'll run you through again

OpenStudy (anonymous):

I Thought we subracted 4 on both sides. But I see that its 3 now ! c: Lol

OpenStudy (bibby):

god my brain is not at full form. We weren't wrong at all hahaha :B 4x=24 x=6 if the amount of roses we have is 6, than the cost is 6*3 (18 dollars) and the amount of daisies we'd have is 9. 18 + 9 = 27

OpenStudy (anonymous):

Okay.. So It would be 18?

OpenStudy (bibby):

yep. the 2nd line is just to make sure it adds up to 27

OpenStudy (anonymous):

WOW! That Was Easy When Someone Actually Explains It! Thank You! :D

OpenStudy (bibby):

Apologies for the confusion. Doing my best :D

OpenStudy (anonymous):

No Its Fine . Thank You !(: